Using Maxwell’s equations in vacuum, derive the wave equation for the z-component
of the electric field vector associated with an electromagnetic wave.
1
Expert's answer
2020-03-09T10:54:32-0400

"\overrightarrow{\nabla}\times\overrightarrow{E}=-\frac{\partial \overrightarrow{B}}{\partial t }"


"\overrightarrow{\nabla}\times\overrightarrow{E}=\mu_0\epsilon_0\frac{\partial \overrightarrow{E}}{\partial t }"


"\overrightarrow{\nabla}\times\overrightarrow{E}(z,t)\overrightarrow{i}=\begin{vmatrix}
 \overrightarrow{i} & \overrightarrow{j} & \overrightarrow{k}\\
 \frac{\partial }{\partial x } & \frac{\partial }{\partial y } & \frac{\partial }{\partial z } \\
\overrightarrow{E}(z,t) &0 & 0
\end{vmatrix}=\frac{\partial E}{\partial z }\overrightarrow{j}"


"\frac{\partial E}{\partial z }=-\frac{\partial B}{\partial t }" (1)


"\overrightarrow{\nabla}\times\overrightarrow{B}(z,t)\overrightarrow{j}=\begin{vmatrix}
 \overrightarrow{i} & \overrightarrow{j} & \overrightarrow{k}\\
 \frac{\partial }{\partial x } & \frac{\partial }{\partial y } & \frac{\partial }{\partial z } \\
0 & \overrightarrow{B}(z,t) & 0
\end{vmatrix}=-\frac{\partial B}{\partial z }\overrightarrow{i}"


"\frac{\partial B}{\partial z }=-\mu_0\epsilon_0\frac{\partial E}{\partial t }"


"\frac{\partial^2 E}{\partial z^2 }=-\frac{\partial}{\partial z }\frac{\partial B}{\partial t }=-\frac{\partial}{\partial t }\frac{\partial B}{\partial z }=-\frac{\partial}{\partial t }(-\mu_0\epsilon_0\frac{\partial E}{\partial t })=\mu_0\epsilon_0\frac{\partial^2 E}{\partial t^2 }"


"\frac{\partial^2 E}{\partial z^2 }=\mu_0\epsilon_0\frac{\partial^2 E}{\partial t^2 }" Answer